OpenStudy (anonymous):

how do you do this? i wanna learn how.. the length of a rectangle is twice the width find the length of the rectangle if the perimeter is 42 cm.. pleaaase explain

OpenStudy (mathteacher1729):

Length = L Width = W "the length of a rectangle is twice the width" as an equation is "L = 2W" The perimeter of a rectangle is 2L + 2W. So we sub L = 2W into (2L + 2W) and obtain (2(2W) + 2W) = 4W + 2W = 6W But the perimeter is given as 42 cm so that means 6W = 42 now solve for W. :)

OpenStudy (anonymous):

Write equations for each aspect of the question. L=Length W=Perimeter L=2W And you know that the perimeter is the sum of the sides of the rectangle, so there are 2 lengths and 2 widths, which gives: 2L+2W=42 Now substitute L=2W into the above equation. 6W=42 W=7 THis means that the length is 14 (L=2W)

OpenStudy (anonymous):

perimeter is \[P=2l+2w\] and you know \[l=2w, P = 42\] so set \[42=2\times 2w+2w\] \[42=6w\] \[\frac{42}{6}=w\] \[7=w\]

OpenStudy (anonymous):

Use the verbal expressions; L=2W Perimeter =2L +2W. Substitute L for 2W and solve!

OpenStudy (anonymous):

would you do 42 divided by 6 t the end?? (thank ya'll!!!!!:Dl)

OpenStudy (anonymous):

at the end*

OpenStudy (mathteacher1729):

Hi Blah blah, yes, you'd divide 42 by 6 at the end. That would give you W. You still have to find L, though. But that is easy, since L is 2W. :)

OpenStudy (anonymous):

tha answer iss...14?:O sweeeeeeeeett.